OpenStudy (anonymous):

PLZ HELP!!!! A launched rocket has an altitude, in meters, given by the polynomial h+vt-4.9t^2, where h is the height, in meters, from which the launch occurs, at velocity v in meters per second, and t is the number of seconds for which the rocket is airborne. If a rocket is launched from the top of a tower 50 meters high with an initial upward speed of 30 meters per second, what will its height be after 2 seconds? h+vt-4.9t^2 Where h = 50 v = 30 t = 2 50+30*2-4.9*2^2 whats the answer?

OpenStudy (turingtest):

You need a calculator?

OpenStudy (anonymous):

lol i have one smarty pants

OpenStudy (turingtest):

50+30*2-4.9*2^2=90.5 there's nothing else to do but use it

OpenStudy (anonymous):

well its saying 90.5 is wrong

OpenStudy (anonymous):

Why does one need a calculator for this computation? 50 + 60 - 4.9 * 4 = 110 - 19.6 = 90.4

OpenStudy (anonymous):

THX GT